OpenStudy (anonymous):

Consider the line that passes through the point (-3, 2) and has a slope of 3. and write using point-slope form

OpenStudy (anonymous):

y-y1=m(x-x1) for (x1,y1) y-2=3(x - (-3) ) y = 3x + 9 + 2 y = 3x +11

OpenStudy (anonymous):

input 2 for y and -3 for x to check the answer 2 = 3(-3) + 11 2 = -9 + 11 2 = 2

OpenStudy (anonymous):

ok, thanks.. how would I put this in slope-intercept form?

OpenStudy (anonymous):

the last line is the point slope solution y = 3x + 11, there are other forms also

OpenStudy (anonymous):

but how would I use this to put in slope-intercept form?

OpenStudy (anonymous):

y-y1=m(x-x1) for (x1,y1)